# Fast dynamical similarity analysis

### v1

@@ Abstract
To understand how neural systems process information, it is often essential to compare one circuit with another, one brain with another, or data with a model. Traditional similarity measures ignore the dynamical processes underlying neural representations. Dynamical similarity methods offer a framework to compare the temporal structure of dynamical systems by embedding their (possibly) nonlinear dynamics into a globally linear space and there computing conjugacy metrics. However, identifying the best embedding and computing these metrics can be computationally slow. Here we introduce fast Dynamical Similarity Analysis (fastDSA), which is computationally far more efficient than previous methods while maintaining their accuracy and robustness. FastDSA introduces two key components that boost efficiency: (1) automatic selection of the effective model order of the Hankel (delay) embedding from the data via a data-driven singular-value threshold that identifies the informative subspace and discards noise to lower computational cost without sacrificing signal, and (2) a novel optimization procedure and objective, which replaces the slow exact orthogonality constraint in finding a minimal distance between dynamics matrices with a lightweight process to keep the search close to the space of orthogonal transformations. We demonstrate that fastDSA is at least an order of magnitude faster than the previous methods. Furthermore, we demonstrate that fastDSA has the properties of its ancestor, including its invariances and sensitivities to system dynamics. FastDSA, therefore, provides a computationally efficient and accurate method for dynamical similarity analysis.
@@
Similarity analysis, Dynamical similarity analysis, Dynamical systems, Neural networks, Neural computations, Computations through dynamics
@@ 1 Main
Neural networks, whether biological or artificial, encode, store, and transform information through their complex dynamics [mante2013context, vyas2020computation, luo2025transitions, pagan2025individual, perich2025neural, chaudhuri2019intrinsic, driscoll2024flexible, jensen2024recurrent, ji2025discovering, genkin2025dynamics, tolmachev2025single, sohn2019bayesian, safavi2024signatures, stringer2024analysis]. Comparing the dynamics of diverse systems is fundamental for uncovering their information processing mechanisms[kriegeskorte2008representational, mante2013context, chaudhuri2016computational, raghu2017svcca, schrimpf2018brain, chaudhuri2019intrinsic, schaeffer2020reverse, vyas2020computation, williams2021generalized, duong2022representational, zeraati2024neural, voges2024decomposing, meilua2024manifold]. Methods that compare dynamics between systems, such as real-world data and models, allow us to gain an algorithmic insight into neural computations [voges2024decomposing] and ultimately a better understanding of the machinery supporting cognitive processes [wang2012neural, zoltowski2020general, cueva2020low, shine2021author, safavi2022multistability, wang2022theory, beiran2023parametric, whyte2024thalamic, safavi2025perceptual, munn2024multiscale, wang2024bifurcation, voigts2025spatial, chandra2025episodic].
Traditional similarity measures primarily focus on the geometry of the data [schrimpf2018brain, kriegeskorte2008representational, raghu2017svcca, williams2021generalized, duong2022representational], or on multivariate statistics, which often struggle to capture the dynamics governing the neural system. Thus, directly comparing the underlying neural dynamics in these systems could be a more faithful approach to compare neural information processing systems, an approach that has recently been gaining currency [eisen2024propofol, ostrow2024delay, guilhot2024dynamical, codol2024brain, bolager2024gradient, huang2024measuring, kujur2025transient, xu2025nn, huang2025inputdsa]. Dynamical Similarity Analysis [DSA, ostrow2023beyond] is a powerful framework designed to compare the governing dynamics in neural networks and beyond. Unlike conventional methods that focus on geometrical and/or statistical similarity, DSA emphasizes the fundamental properties of systems’ dynamics by leveraging insights from the theory of dynamical systems [thompson2002nonlinear, strogatz2024]. DSA provides a principled way to measure how much two systems exhibit equivalent dynamics, even when their trajectories in the state-space appear geometrically distinct.
Unlike traditional methods [schrimpf2018brain, kriegeskorte2008representational, raghu2017svcca, williams2021generalized, duong2022representational], DSA targets the similarity of dynamical systems by estimating their topological conjugacy. It does so by embedding nonlinear dynamics into a globally linear space using Koopman theory [snyder2021koopman]. It then directly compares flow fields, which represent the evolution of the system’s dynamics over time by aligning them under optimal orthogonal transformations to assess their similarity. This novel approach allows DSA to differentiate between systems with (possibly) similar geometries but fundamentally different dynamics. It also identifies dynamical similarities between systems with distinct spatial configurations and/or architectures.
The analysis of dynamical similarity provides a crucial tool for validating computational models of information processing. For instance, let us consider the example of a Recurrent Neural Network (RNN) trained on a center-out reaching task, against activity recorded from the primary motor cortex (M1) of a monkey performing the same task [codol2024brain]. Traditional comparison methods, which focus on the geometry (or a static picture) of the high-dimensional neural trajectories, often fail because arbitrary scaling or rotation between the biological and artificial state-spaces can lead to a false conclusion of low similarity, even if the underlying computations are identical. Dynamic similarity analysis can overcome such caveats by comparing the governing dynamics, exploiting the flow field of the system’s dynamics (which in a way, specifies the instantaneous transformation rule). By establishing whether a single linear map can align the dynamics of the RNN with the dynamics of the M1 neuronal populations, dynamic similarity analysis yields a high similarity score if the RNN is implementing the same computational mechanism as the brain, thus moving validation beyond task performance and to a mechanistic understanding of neural computations.
In this study, we focus on improving the computational efficiency of DSA to broaden its usability. As discussed in detail in Methods section, one of the key computational bottlenecks in DSA, among other factors, is the optimization process required to align vector fields. Here, we suggest incorporating hybrid approaches for optimization and extending DSA, which can considerably improve the computational efficiency. Thus, we introduce fastDSA, an optimized approach, designed to accelerate DSA computations while maintaining the accuracy and robustness of dynamic similarity estimation.
@@ 2 Results
FastDSA is an efficient and accurate method for estimating dynamic (dis)similarity between neural networks, neuronal circuits, or data with a model. In fastDSA, we integrate concepts from random matrix theory to identify a low-rank approximation of the system’s dynamics, and develop new optimization procedures for aligning vector fields. We introduce and compare three variants of our fastDSA (each based on a different optimization procedure), and show that all variants preserve the accuracy and mathematical properties of its predecessor [ostrow2023beyond] while substantially enhancing computational efficiency. We demonstrate the capacity of the family of fastDSA methods with several numerical experiments with systematic manipulation of governing dynamics in neural networks.
2.1 Accelerating dynamical similarity analysis
The core architecture of fastDSA is based on the original DSA [ostrow2023beyond]; therefore, we first provide the key formulation of the original DSA [see, ostrow2023beyond, for mathematical details of the formulation] and then the novel developments of fastDSA. Let $x$ and $y$ be two dynamical systems:
@@ 2.1 Accelerating dynamical similarity analysis
Then, data from each system can be structured into Hankel matrices ($H_{x}$, $H_{y}$). The delay embedding effectively lifts the nonlinear dynamics into a higher-dimensional space where, according to Koopman theory (in the limit of infinite-dimensional embeddings), they can be approximated as linear [arbabi2017ergodic, brunton2021modernkoopmantheorydynamical]:
@@
where $H_{x}^{T}=U_{x}\Sigma_{x}V_{x}^{T}$ is the singular value decomposition (SVD) of the transposed Hankel matrix, $H$; $\Sigma$, and $V$ are the standard SVD components; $r$ denotes the chosen truncation rank; $V_{x}[:,1:r]$ consists of the first $r$ columns of $V_{x}$; $A_{x}$ is the matrix that provides the approximate linear dynamics in the reduced subspace, and $V^{\prime}_{x,r}$ represents the next predicted step in this coordinate space. The resulting data matrices represent the approximated linear dynamics. This numerical methodology, also known as the Dynamic Mode Decomposition (DMD), serves as a practical method for analyzing complex nonlinear dynamics through a linear approximation [brunton2017chaos]. The DMD matrix $A$ is typically decomposed into eigenvectors and eigenvalues, which capture core dynamic properties of the data. Theoretically, the DMD seeks a finite approximate of the Koopman operator [arbabi2017ergodic, brunton2021modernkoopmantheorydynamical] that linearizes the evolution of functions on the state space. The DMD identifies a linear operator that can estimate the system’s dynamics in the next time step $t+\Delta t$, $X_{(t+\Delta t)}=AX_{(t)}$. For the lifting and linearization step, the original DSA employed delay embeddings combined with DMD. However, the general framework of DSA is not limited to this specific choice of embedding. Other approaches (such as kernel methods, neural network encoders, or alternative Koopman operator approximations) can likewise be used to map nonlinear dynamics into a space where linear comparisons are meaningful. On the next step, the linear dynamics approximations $A_{x},A_{y}$, are compared using a modified Procrustes analysis tailored specifically for vector fields (Procrustes Analysis over Vector Fields, PAVF):
@@
To computationally accelerates the originally developed DSA [ostrow2023beyond], we introduce two major changes: (1) Calculating the approximate optimal rank automatically (using methods based on random matrix theory). (2) Incorporating novel optimization constraints for PAVF to reduce computational overhead (see Figure 1 for a schematic illustration of the family of fastDSA methods and their comparison; and section Computational foundation of fastDSA and Supp. Figure 1 for further elaboration).
Figure 1: Schematic overview of methods for estimating dynamic (dis)similarity: DSA, family of fastDSA methods, and kwDSA: (Top) Depict the overall pipeline common to all methods. In particular, the similarity computation is decomposed into three consecutive steps: lifting to a higher-dimensional representation (purple), $(x_{1},x_{2},\ldots,x_{n})\;\longmapsto\;(\tilde{x}_{1},\tilde{x}_{2},\ldots,\tilde{x}_{m})$, where $m>n$, linear approximation of the dynamics based on the Koopman theory (green), and optimization of the similarity transform (brown). (Bottom) The chart summarizes how each method (DSA, the three fastDSA variants, and kwDSA) instantiates these three stages. In summary, all three fastDSA variants employ automatic rank detection in the linear approximation step. For kwDSA there is an implicit lifting to higher dimension through a kernel function [see section Sensitivity to fine changes in governing dynamics and bernhard2004, bernhard2001, for more details]. Furthermore, for kwDSA, the automatic rank selection can also be applied. At the optimization stage, however, the methods differ substantially, including the three fastDSA variants, which leads to distinct computational and accuracy profiles across algorithms (see section Assessing efficacy and reliability of fastDSA for more details).
We develop three novel optimizers for the alignment problem in DSA that remain close to the original formulation, in terms of mathematical properties, while improving computational efficiency. These provide alternative ways to find the similarity transform for the orthogonal group $O(n)$ (or $SO(n)$). Given two matrices $A_{x}$ and $A_{y}$, we seek a matrix $C$ such that
@@
In the original DSA, this constraint is reimposed after every update step, which slows convergence. We introduce three variants to remedy this bottleneck: (1) fastDSA with a regularized optimization (including a penalty term in the optimization procedure (we will refer to as RO-fastDSA). (2) fastDSA that is optimized based on-manifold Riemannian method [ablin2022fast] that we will refer to as Rim-fastDSA). (3) We use the Landing algorithm [ablin2022fast] to optimize for the alignment problem of DSA (we will refer to it as Land-fastDSA). See section Computational foundation of fastDSA for the details of each optimization variant. In the next sections, we demonstrate the practical implications of these novel optimizations along with other key components of fastDSA.
@@ 2.2 Automatic estimation of optimal rank of dynamical systems
One of the main challenges of the DMD is to determine the appropriate rank for the SVD that is used in the linear representation of the system (also, see, Eq. 2). Selecting a rank that is too high can introduce spurious modes into the comparison, while a rank that is too low may cause the loss of essential information. Furthermore, using an optimal rank can also potentially act as denoising, which facilitates estimation, given that noise components typically correspond to smaller singular values. Several methods have been proposed for automated and statistically-grounded rank selection [nadlerDistributionRatioLargest2011, kritchmanNonParametricDetectionNumber2009, kritchmanDeterminingNumberComponents2008, veraartDenoisingDiffusionMRI2016, safaviUnivariateMultivariateCoupling2021, safaviUncoveringOrganizationNeural2023]. One notable approach is the Optimal Singular Value Hard Thresholding [SVHT, gavish2014optimal], which forms the basis of our method for estimating the optimal rank of dynamical systems and has a solid statistical foundation based on random matrix theory [plerouRandomMatrixApproach2002, paulRandomMatrixTheory2014, leeRandomMatrixTheory2014, izenmanRandomMatrixTheory, karouiSpectrumEstimationLarge2008].
@@
where $x_{\text{med}}$ is the median singular value of $X$, and $\omega(\beta)$ is a precomputed constant [gavish2014optimal].
@@
with standard chaotic parameters set as $\sigma=10$, $\rho=28$, and $\beta=\frac{8}{3}$. The true latent state at any time is represented by the three-dimensional vector $(x,y,z)$, confirming an intrinsic rank of 3 (Figure 2a, top-left). Thus, we simulated Lorenz trajectories by initializing the system at $x_{0},y_{0},z_{0}]=[0,1,1.05]$ and numerically integrating over time. This process generates a multivariate timeseries with a known low-dimensional latent dynamics (with three dimensions and length $T$, represents the number of time steps). Next, to simulate observational data with higher dimensionality, a random projection matrix $W\in\mathbb{R}^{3\times D}$ was constructed, where $D$ is the number of observed dimensions (in this case 128). Application of the projector operator $W$ leads to a new higher-dimensional ($D$-dimensional) data matrix that retained an approximate intrinsic rank of 3, reflecting its low-dimensional origin [similar to, kapoor2024latent, ; see Figure 2a, top-right]. We then demonstrate the effectiveness of our method for rank detection by testing it on dynamics with known rank.
To further assess SVHT as a method for identifying optimal rank, we compared its rank selection accuracy against traditional information criteria such as the Akaike Information Criterion (AIC), Bayesian Information Criterion (BIC), and the corrected AIC [AICc, hurvich1993corrected, see Figure 2a, bottom-left]. The conventional formulations of AIC and BIC were used for estimation here:
@@
We also demonstrate the SVHT robustness for rank detection in the presence of noise. This capacity stems from its foundation in random matrix theory that exploit the known reflection of noise statistics in the spectrum of singular values [nadlerDistributionRatioLargest2011, kritchmanNonParametricDetectionNumber2009, kritchmanDeterminingNumberComponents2008, veraartDenoisingDiffusionMRI2016, safaviUnivariateMultivariateCoupling2021, safaviUncoveringOrganizationNeural2023]. Thus, we used varying levels of noise to the projected data to assess the robustness of the SVHT-based method for determining the optimal rank. SVHT consistently identified the correct rank of the latent dynamics (rank 3) in moderately noisy conditions (Figure 2a, bottom-right).
@@
Figure 2: Demonstration of automatic rank reduction with SVHT: (a) (top-left) Synthetic Lorenz attractor. (top-right) Depiction of the projected Lorenz attractor timeseries into a higher-dimensional (128-dimensional) observation space [similar to, kapoor2024latent]. X-axis of the heatmap indicates time steps, and the y-axis indicates the dimensions. (bottom-left) Comparison of optimal rank detection by SVHT (brown broken line) against traditional model selection criteria (AIC, red; AICc, purple; BIC, green). Due to the overlap, they are not all fully visible. SVHT accurately identifies the rank of underlying latent dynamics (rank 3) that precisely overlap with the ground truth (GT, vertical continuous blue line). AIC, AICc (corrected AIC for small samples), and BIC after the knee point on $r=3$, exhibit a monotonic decrease in their values with increasing rank $r$. (bottom-right) Stability of SVHT-detected optimal rank under varying levels of additive noise. (b) Schematics of the process of injecting noise to synthetic Lorenz attractor data with isotropic noise (left) or non-linear noise (right), both were generated with variance $\sigma=5$, and (bottom) DMD reconstructed data using the optimal rank detected by SVHT. (c-d) Mean Squared Error (MSE), on left, and AIC values, on right, for DMD-based reconstruction across varying levels of (c) isotropic and (d) nonlinear noise. (e) (top) MSE and (bottom) AIC of DMD-based reconstruction under nonlinear noise, and the SVHT-selected optimal rank (brown broken vertical line) for different values of SNR. Darker colors correspond to higher SNR values (lower noise levels), and brighter colors to lower SNR values (higher noise levels). SNR values are noted on top of individual columns.
2.3 Assessing novel optimizers for dynamic similarity analysis
We first demonstrate that all fastDSA optimizers we introduce for the alignment problem discussed earlier are superior to the one used in DSA [ostrow2023beyond]. Thus, we compare and characterize the variability across different types of optimizers. To conduct this comparison and characterize the differences, we first designed an experiment to quantify speed–accuracy trade-offs under closely matched conditions. The testbed consisted of generating random square Gaussian matrices with size $n\in\{2,4,8,16,32,64,128,256\}$. For each $n$, we first generated synthetic symmetric positive definite (SPD) matrices with controlled spectral properties. Given a matrix size $n$ and a minimum eigenvalue threshold $\lambda_{\min}>0$, we drew a random matrix $A\in\mathbb{R}^{n\times n}$ with independent standard normal entries and symmetrized it as

### v2

@@ Abstract
To understand how nonlinear dynamical systems process information (e.g., artificial neural networks and neural circuits), it is essential to compare the underlying dynamics of diverse systems at scales (e.g., a large pool of neural networks with diverse architectures and large-scale recordings from neural circuits). Although a substantial body of similarity metrics has been developed, existing approaches remain inadequate for large-scale comparisons of systems. On one hand, geometric similarity methods are computationally efficient and allow large-scale comparisons; however, they are inaccurate as they do not capture the governing dynamics, which is essential to understand these nonlinear dynamical systems and the computations they implement. On the other hand, conventional methods to compare the similarity of dynamics are faithful to governing dynamics but often impractical due to their computational cost. Here, we bring the best of both worlds, the computational efficiency of geometric tools and the reliability of dynamic similarity methods, enabling us to compare diverse nonlinear dynamical systems at large scales. We introduce fast dynamical similarity analysis (fastDSA) as a computationally efficient and accurate (dis)similarity metric between two nonlinear dynamical systems. FastDSA is built based on modern computational methods: methods from random matrix theory for the detection of the optimal rank of a dynamical system, developing multiple novel optimization pipelines to align the flow field of dynamical systems, and powerful Koopman embeddings. We demonstrate that across multiple benchmark nonlinear systems and recurrent network models, fastDSA is insensitive to arbitrary coordinate choices, but sensitive to real differences in the systems’ dynamics (how the systems evolve over time), which geometric methods may miss and are computationally expensive with traditional dynamical methods. To the best of our knowledge, fastDSA is the fastest yet accurate method to compare the (dis)similarity of nonlinear dynamical systems. FastDSA allows large-scale and statistical analyses of a wide range of nonlinear dynamical systems, consequently, extending the practical scope of dynamical similarity analysis.
@@
Nonlinear dynamics, Dynamical systems, Similarity analysis, Dynamical systems analysis at scale, Neural networks, Neural computations, Computations through dynamics
@@ 1 Main
Nonlinear dynamical systems serve as a robust foundation for diverse information processing architectures [81, 39, 20, 55, 88]. In particular, neural networks, whether biological or artificial, encode, store, and transform information through their complex dynamics [50, 55, 88, 48, 59, 61, 15, 20, 32, 33, 24, 84, 78, 68, 79]. Comparing the dynamics of diverse systems is fundamental for uncovering their information processing mechanisms [41, 50, 14, 64, 75, 15, 72, 88, 97, 21, 101, 86, 51]. Methods that compare dynamics between systems, such as real-world data and models, allow us to gain an algorithmic insight into neural computations [86] and ultimately a better understanding of the machinery supporting cognitive computations [94, 104, 19, 76, 69, 95, 7, 96, 53, 93, 87, 13]. Thus, tools to compare the dynamics of diverse neural networks at scales are essential to understand how they process information.
Although numerous similarity metrics have been proposed to compare neural networks, they are not practical for studying the dynamics of neural networks with diverse characteristics on large scales — which is crucial to understand their dynamical principles serving information processing [20, 55, 88]. A large subclass of similarity measures primarily focuses on the geometry of the data, or on multivariate statistics [75, 41, 64, 97, 21, 5]. While these methods are computationally efficient, they often struggle to capture the dynamics governing the neural system, which is essential to understand how they realize computations [57, 28]. Thus, comparing the underlying neural dynamics in these systems was suggested as a more faithful approach to compare neural information processing systems, an approach that has recently been gaining currency [57, 22, 56, 47, 16, 26, 17, 9, 29, 44, 100, 28]. Unlike geometrical methods that focus on geometrical and/or statistical similarities, methods for dynamic similarity analysis [22, 56, 26, 17, 9, 29, 44, 100, 28] emphasize the fundamental properties of systems’ dynamics by leveraging insights from the theory of dynamical systems [83, 80]. This new class of (dynamic) similarity metrics is conceptually appealing; however, their high computational cost makes them impractical for large-scale analysis. Specifically, it is too expensive to deploy across extensive pools of networks featuring diverse architectures, input regimes, and learning curricula. The ideal solution combines the best of both approaches: the computational efficiency of geometrical methods and the accuracy of dynamic similarity metrics.
In this study, we introduce fast dynamical similarity analysis (fastDSA) that has the advantages of both worlds: accuracy and mathematical properties of methods based on dynamic similarity analysis and the computational efficiency of geometrical methods. FastDSA uses the principles of Koopman theory to capture the most essential aspects of nonlinear dynamical systems [similar to, 22, 56, 26, 17, 29, 44, 100, 28] and targets the similarity of dynamical systems by estimating their topological conjugacy. It does so by embedding nonlinear dynamics into a globally linear space using Koopman theory [77]. To make this comparison practical at scale, fastDSA introduces automated model-order (rank) selection for Hankel–Koopman embeddings and efficient optimization schemes for near-orthogonal alignment, substantially reducing the computational bottlenecks of operator-alignment dynamical similarity. As a result, it can distinguish systems with similar state-space geometry but different dynamics, and it can identify dynamical similarity across systems with distinct spatial configurations and/or architectures.
Through a set of diverse analyses, on simulated nonlinear dynamical systems and recurrent neural networks, we demonstrate that fastDSA successfully combines the strength of both worlds (computational efficacy and reliable dynamic similarity assessment) for comparing the (dis)similarity of nonlinear dynamical systems. First, using a ring attractor network, we show that fastDSA maintains strict invariance to topology-preserving geometric deformations. Second, by parametrically transforming the ring into a line attractor, we confirm that fastDSA accurately detects genuine topological shifts in the governing dynamics. Finally, we compare fastDSA to fast baseline similarity measures. This analysis tests two practical desiderata: robustness to noise and sensitivity to fine-grained differences in dynamical mechanisms. Notably, we demonstrate in all scenarios we evaluate fastDSA was computationally more efficient than conventional dynamical similarity metrics and more accurate than geometrical methods. Altogether, we demonstrate fastDSA is a computationally efficient, mathematically solid, and accurate framework to compare dynamics across nonlinear dynamical systems (in particular neural networks). Most importantly, fastDSA enables large-scale analysis of nonlinear dynamical systems across models, training conditions, and experimental datasets.
@@ 2 Results
FastDSA is an efficient and accurate method for estimating dynamic (dis)similarity between neural networks, neuronal circuits, or data with a model. In fastDSA, we integrate concepts from random matrix theory to identify a low-rank approximation of the system’s dynamics, and develop new optimization procedures for aligning vector fields. We introduce and compare three variants of our fastDSA (each based on a different optimization procedure), and show that all variants preserve the accuracy and mathematical properties of conventional dynamical similarity methods [57] while substantially enhancing computational efficiency. We demonstrate the capacity of the family of fastDSA methods with several numerical experiments with systematic manipulation of the governing dynamics in neural networks.
2.1 A scalable Koopman operator alignment for dynamical similarity
Our goal is to quantify dynamical similarity by comparing governing dynamics rather than the geometry of data (e.g., statistical similarities). We do so by constructing Koopman-based linear representations of each system and then measuring similarity through an operator-alignment: if two systems implement equivalent dynamics up to a change of coordinates, their induced operators can be aligned by an (approximately) orthogonal similarity transform, and if different, they will be proportionally further away from an orthogonal similarity transform. We formalize as follows. Let $x$ and $y$ be two dynamical systems:
@@ 2.1 A scalable Koopman operator alignment for dynamical similarity
Then, data from each system can be structured into Hankel matrices ($H_{x}$, $H_{y}$). The delay embedding effectively lifts the nonlinear dynamics into a higher-dimensional space where, according to Koopman theory (in the limit of infinite-dimensional embeddings), they can be approximated as linear [2, 11]:
@@
where $H_{x}^{T}=U_{x}\Sigma_{x}V_{x}^{T}$ is the singular value decomposition (SVD) of the transposed Hankel matrix, $H$; $\Sigma$, and $V$ are the standard SVD components; $r$ denotes the chosen truncation rank; $V_{x}[:,1:r]$ consists of the first $r$ columns of $V_{x}$; $A_{x}$ is the matrix that provides the approximate linear dynamics in the reduced subspace, and $V^{\prime}_{x,r}$ represents the next predicted step in this coordinate space. The resulting data matrices represent the approximated linear dynamics. This numerical methodology, also known as the Dynamic Mode Decomposition (DMD), serves as a practical method for analyzing complex nonlinear dynamics through a linear approximation [10]. The DMD matrix $A$ is typically decomposed into eigenvectors and eigenvalues, which capture core dynamic properties of the data. Theoretically, the DMD seeks a finite approximate of the Koopman operator [2, 11] that linearize the evolution of functions on the state space. The DMD identifies a linear operator that can estimate the system’s dynamics in the next time step $t+\Delta t$, $X_{(t+\Delta t)}=AX_{(t)}$. In previous studies [2, 10, 57, 36] the lifting step has often been instantiated using delay-coordinate (Hankel) embeddings followed by DMD/Hankel-DMD (HAVOK) to obtain a finite-dimensional Koopman approximation. Here, we adopt the same practical Koopman-embedding strategy; however, we emphasize that the broader operator-alignment view [that is not limited to this choice, kernel methods, neural encoders, or alternative Koopman approximations can likewise map nonlinear dynamics into a representation where linear operator comparisons are meaningful; also see, 58]. On the next step, the linear dynamics approximations $A_{x},A_{y}$, are compared using a modified Procrustes analysis tailored specifically for vector fields (Procrustes Analysis over Vector Fields, PAVF):
@@
FastDSA builds on well-established operator-alignment approaches [e.g., 57, 16, 65, 102, 36], but removes two bottlenecks that have limited existing methods to modest problem sizes: (1) Calculating the approximate optimal rank automatically (using methods based on random matrix theory). (2) Incorporating novel optimization constraints for PAVF to reduce computational overhead (see Figure 1 for a schematic illustration of the family of fastDSAmethods and their comparison; and section Computational foundation of fastDSA and Supp. Figure 1 for further elaboration).
Figure 1: Schematic overview of methods for estimating dynamic (dis)similarity: DSA, family of fastDSA methods, and kwDSA: (Top) Depict the overall pipeline common to all methods. In particular, the similarity computation is decomposed into three consecutive steps: lifting to a higher-dimensional representation (purple), $(x_{1},x_{2},\ldots,x_{n})\;\longmapsto\;(\tilde{x}_{1},\tilde{x}_{2},\ldots,\tilde{x}_{m})$, where $m>n$, linear approximation of the dynamics based on the Koopman theory (green), and optimization of the similarity transform (brown). (Bottom) The chart summarizes how each method (DSA, the three fastDSA variants, and kwDSA) instantiates these three stages. In summary, all three fastDSA variants employ automatic rank detection in the linear approximation step. For kwDSA there is an implicit lifting to higher dimension through a kernel function [see section Sensitivity to fine changes in governing dynamics and 74, 73, for more details]. Furthermore, for kwDSA, the automatic rank selection can also be applied. At the optimization stage, however, the methods differ substantially, including the three fastDSA variants, which leads to distinct computational and accuracy profiles across algorithms (see section Assessing efficacy and reliability of fastDSA for more details).
Building on prior operator-alignment formulations, we propose three alignment optimizers that retain the same objective and invariance guarantees while making the constrained alignment problem computationally tractable in large-sample settings. These provide alternative ways to find the similarity transform for the orthogonal group $O(n)$ (or $SO(n)$). Given two matrices $A_{x}$ and $A_{y}$, we seek a matrix $C$ such that
@@
In prior operator-alignment formulations [e.g., in 57], the orthogonality constraint is typically reimposed after each update step, which can dominate runtime and hinder scalability. FastDSA addresses this bottleneck by supporting three alignment solvers that optimize the same operator-alignment objective while handling near-orthogonality more efficiently: (1) fastDSA with a regularized optimization (including a penalty term in the optimization procedure (we will refer to as RO-fastDSA). (2) fastDSA that is optimized based on-manifold Riemannian method [1] that we will refer to as Rim-fastDSA). (3) We use the Landing algorithm [1] to optimize for the alignment problem of DSA (we will refer to it as Land-fastDSA). See section Computational foundation of fastDSA for the details of each optimization variant. In the next sections, we demonstrate the practical implications of these novel optimizations along with other key components of fastDSA.
@@ 2.2 Automatic estimation of optimal rank of dynamical systems
One of the main challenges of the DMD is to determine the appropriate rank for the SVD that is used in the linear representation of the system (also, see, Eq. 2). Selecting a rank that is too high can introduce spurious modes into the comparison, while a rank that is too low may cause the loss of essential information. Furthermore, using an optimal rank can also potentially act as denoising, which facilitates estimation, given that noise components typically correspond to smaller singular values. Several methods have been proposed for automated and statistically-grounded rank selection [54, 43, 42, 85, 70, 71]. One notable approach is the Optimal Singular Value Hard Thresholding [SVHT, 23], which forms the basis of our method for estimating the optimal rank of dynamical systems and has a solid statistical foundation based on random matrix theory [63, 60, 46, 31, 38].
@@
where $x_{\text{med}}$ is the median singular value of $X$, and $\omega(\beta)$ is a precomputed constant [23].
@@
with standard chaotic parameters set as $\sigma=10$, $\rho=28$, and $\beta=\frac{8}{3}$. The true latent state at any time is represented by the three-dimensional vector $(x,y,z)$, confirming an intrinsic rank of 3 (Figure 2a, top-left). Thus, we simulated Lorenz trajectories by initializing the system at $x_{0},y_{0},z_{0}]=[0,1,1.05]$ and numerically integrating over time. This process generates a multivariate timeseries with a known low-dimensional latent dynamics (with three dimensions and length $T$, representing the number of time steps). Next, to simulate observational data with higher dimensionality, a random projection matrix $W\in\mathbb{R}^{3\times D}$ was constructed, where $D$ is the number of observed dimensions (in this case 128). Application of the projector operator $W$ leads to a new higher-dimensional ($D$-dimensional) data matrix that retained an approximate intrinsic rank of 3, reflecting its low-dimensional origin [similar to, 37, ; see Figure 2a, top-right]. We then demonstrate the effectiveness of our method for rank detection by testing it on dynamics with known rank.
To further assess SVHT as a method for identifying optimal rank, we compared its rank selection accuracy against traditional information criteria such as the Akaike Information Criterion (AIC), Bayesian Information Criterion (BIC), and the corrected AIC [AICc, 30, see Figure 2a, bottom-left]. The conventional formulations of AIC and BIC were used for estimation here:
@@
We also demonstrate the SVHT robustness for rank detection in the presence of noise. This capacity stems from its foundation in random matrix theory that exploit the known reflection of noise statistics in the spectrum of singular values [54, 43, 42, 85, 70, 71]. Thus, we used varying levels of noise to the projected data to assess the robustness of the SVHT-based method for determining the optimal rank. SVHT consistently identified the correct rank of the latent dynamics (rank 3) in moderately noisy conditions (Figure 2a, bottom-right).
@@
Figure 2: Demonstration of automatic rank reduction with SVHT: (a) (top-left) Synthetic Lorenz attractor. (top-right) Depiction of the projected Lorenz attractor timeseries into a higher-dimensional (128-dimensional) observation space [similar to, 37]. X-axis of the heatmap indicates time steps, and the y-axis indicates the dimensions. (bottom-left) Comparison of optimal rank detection by SVHT (brown broken line) against traditional model selection criteria (AIC, red; AICc, purple; BIC, green). Due to the overlap, they are not all fully visible. SVHT accurately identifies the rank of underlying latent dynamics (rank 3) that precisely overlap with the ground truth (GT, vertical continuous blue line). AIC, AICc (corrected AIC for small samples), and BIC after the knee point on $r=3$, exhibit a monotonic decrease in their values with increasing rank $r$. (bottom-right) Stability of SVHT-detected optimal rank under varying levels of additive noise. (b) Schematics of the process of injecting noise to synthetic Lorenz attractor data with isotropic noise (left) or non-linear noise (right), both were generated with variance $\sigma=5$, and (bottom) DMD reconstructed data using the optimal rank detected by SVHT. (c-d) Mean Squared Error (MSE), on left, and AIC values, on right, for DMD-based reconstruction across varying levels of (c) isotropic and (d) nonlinear noise. (e) (top) MSE and (bottom) AIC of DMD-based reconstruction under nonlinear noise, and the SVHT-selected optimal rank (brown broken vertical line) for different values of SNR. Darker colors correspond to higher SNR values (lower noise levels), and brighter colors to lower SNR values (higher noise levels). SNR values are noted on top of individual columns.
2.3 Assessing novel optimizers for dynamic alignment comparison
We first demonstrate that all fastDSA’s optimizers we introduce for the alignment problem discussed earlier are superior to previous methods for dynamic similarity analysis [57]. Thus, we compare and characterize the variability across different types of optimizers. To conduct this comparison and characterize the differences, we first designed an experiment to quantify speed–accuracy trade-offs under closely matched conditions. The testbed consisted of generating random square Gaussian matrices with size $n\in\{2,4,8,16,32,64,128,256\}$. For each $n$, we first generated synthetic symmetric positive definite (SPD) matrices with controlled spectral properties. Given a matrix size $n$ and a minimum eigenvalue threshold $\lambda_{\min}>0$, we drew a random matrix $A\in\mathbb{R}^{n\times n}$ with independent standard normal entries and symmetrized it as